# Switching the Loss Reduces the Cost in Offline Reinforcement Learning

### v2

@@ 5.3 Asterix and Seaquest
We evaluate the deep RL variants of FQI-log and FQI-sq on the Atari 2600 games Asterix and Seaquest (Bellemare et al., 2013), and use the distributional RL algorithm C51 as an additional baseline. We adopt the data and experimental setup of Agarwal et al. (2020). The data consist of five batch datasets for each game, which were collected from independent training runs of a DQN agent (Mnih et al., 2015). Specifically, each batch dataset contains every fourth frame from 200 million frames of training; a frame skip of four and sticky actions (Machado et al., 2018) were used, whereby all actions were repeated four times consecutively and an agent randomly repeated its previous action with probability 0.25.
When the function class $\mathcal{F}$ of FQI-sq is given by a parameterized deep neural network, the algorithm is called DQN. To adapt FQI-log to the deep RL setting, we must switch the training loss from $\ell_{\text{sq}}$ to $\ell_{\log}$, and add a sigmoid activation layer to squash the output range to [0,1]. We henceforth refer to these algorithms as DQN-sq and DQN-log, respectively.
@@
    C51 replaces the Bellman optimality operator $\mathcal{T}$ with a modified “distributional Bellman operator” for the updates.
For our experiments we clip the targets of DQN-log in the same way as C51, and we set $V_{\min}=0$ and $V_{\max}=10$. Since the sigmoid activation of DQN-log is a specialization of the softmax activation to the binary case, DQN-log implements the changes S.3, S.2, and S.4 to the standard form of DQN-sq. Clipping the targets introduces a bias which we correct for by similarly clipping the targets of DQN-sq. Thus our benchmark results for DQN-sq include the change S.4. Clipping the targets of DQN-sq is novel to this work and improves performance, yielding a stronger baseline. We include a comparison with the traditional unclipped version of DQN-sq in Appendix C.
Agarwal et al. (2020) found that C51 significantly outperforms DQN-sq on the games Asterix and Seaquest. We selected these games to test our hypothesis that the improved performance of C51 can be, in part, attributed to the choice of loss function. In our implementations of DQN-log and DQN-sq, we use the same hyperparameters reported by Agarwal et al. (2020). Fig. 2 shows the undiscounted return as a function of the number of training epochs. On Seaquest, DQN-log outperforms DQN-sq and is able to match the performance of C51. In Asterix, DQN-log performs similarly to DQN-sq and both achieve lower return than C51.
5.4 Experimental Findings
Our experimental results suggest that log-loss offers improvements over squared loss, as demonstrated by the results of FQI-log on Mountain Car and DQN-log on Seaquest. Furthermore, the deep RL experiments show that DQN-log can sometimes even perform similarly to the strong distributional baseline of C51. These improvements are gained by simply modifying a few lines of code, whereas C51’s implementation is much more complex.
Figure 2: Learning curves on Asterix and Seaquest. The result are averaged over 5 datasets. The shaded regions represent one standard error of the mean. One epoch contains 100k updates.
@@ 6 Related Works
Wang et al. (2023) obtain small-cost bounds for finite-horizon batch RL problems under the distributional Bellman completeness assumption, which is stronger than the one made in Assumption 4.5. And, Wang et al., 2024 refines the bounds of Wang et al., 2023, showing second-order bounds (depending on variance) for the same algorithm proposed in the latter. They attribute their small-cost bound to the use of the distributional Bellman operator (Bellemare et al., 2023). However, their proof techniques only make use of pessimism (Buckman et al., 2021; Jin et al., 2021) and log-loss in achieving their small-cost bound. The use of pessimism is necessary for their proof in order to control the errors accumulated by use of the distribution Bellman operator during value iteration. We improve upon their work by proposing an efficient algorithm for batch RL that enjoys a similar small-cost bound without use of the distributional Bellman operator or pessimism under a weaker completeness assumption.
Jin et al. (2020a) and Wagenmaker et al. (2022) obtain regret bounds that scale with the value of the optimal policy. However in their setting, the goal is to maximize reward. Therefore, their bounds only improve upon the previous bounds (Azar et al., 2017; Yang and Wang, 2019; Jin et al., 2020b) when the optimal policy accumulates very little reward. These bounds are somewhat vacuous as they only imply that regret is low when the value of the initial policy is already close to the value of the optimal policy, both of which are close to zero. Small-cost bounds give the same rates as small-return bounds, however, they are more attractive as the cost of the initial policy can be high while the cost of the optimal policy can be low. Small-cost bounds for tabular MDPs where given by Lee et al. (2020) and for linear quadratic regulators (LQR) by Kakade et al. (2020). To our knowledge we give the first small-cost bound for batch RL under the standard completeness assumption (Antos et al., 2007; Munos and Szepesvári, 2008).
@@
Several works get small-cost bounds in contextual bandits (Allen-Zhu et al., 2018; Foster and Krishnamurthy, 2021; Olkhovskaya et al., 2023). Foster and Krishnamurthy (2021) show that for cost sensitive classification, a batch variant of contextual bandits, learning a value function via log-loss enjoys a small-cost bound, see Theorem 3. Whereas learning a value function via squared loss fails to achieve a small-cost guarantee, see Theorem 2. Our main result, Theorem 4.7 can be viewed as the sequel to Theorem 3 of Foster and Krishnamurthy (2021).
Abeille et al. (2021) show that for stochastic logistic bandits, where the costs/rewards are Bernoulli, the bounds scale with the variance of the optimal arm. This is simultaneously a small-cost and small-return bound. They achieve this bound by computing an optimistic estimate of the Bernoulli maximum likelihood estimator (MLE). Janz et al. (2023) extend this result and shows that for any “self-concordant” random variable computing the MLE and being optimistic gives a bound that scales with the variance of the optimal arm.
@@
The theoretical literature on batch RL has largely been focused on proving sample efficiency rates. Chen and Jiang (2019) proved that FQI-sq gets a rate optimal bound of $1/\sqrt{n}$ when realizability, concentrability and completeness hold. Foster et al. (2021) then show that if one assumes concentrability then completeness is a necessary assumption for sample efficient batch RL. Xie and Jiang (2021) prove that if one uses the stronger notation of concentrability from Munos (2003) then sample efficient batch RL is possible with only the added assumption of realizability, i.e. the optimal value function lies in the function class. To our knowledge the previous theoretical works on batch RL has largely focused on algorithms optimized via squared loss (Antos et al., 2007; Farahmand, 2011; Pires and Szepesvári, 2012; Chen and Jiang, 2019; Xie and Jiang, 2021).
@@ 7 Conclusions
We showed that in batch RL the loss function genuinely matters. We proved that FQI-log is more sample efficient on MDPs where the optimal cost $\bar{v}^{\star}$ is small. FQI-log is the first efficient batch RL algorithm to get a small-cost bound.
We believe our result holds generally and can be extended to any batch RL setting where the squared loss has bounded error. An intriguing extension would be deriving small-cost bounds in batch RL with only realizability (and a stronger concentrability assumption). Xie and Jiang (2021) show that with only a realizable function class squared loss gets bounded error and we conjecture a small-cost bound could be shown by adapting the analysis detailed in our work to their setting.
Wagenmaker et al. (2022) get small-return bounds for online RL in linear MDPs (Jin et al., 2020b). Another interesting avenue for future work would be getting small-cost bounds in online RL with linear function approximation. Finally, our Mountain Car experiments indicate log-loss might perform well in goal-oriented MDPs (Bertsekas, 1995). In goal-oriented MDPs the agent is tasked with reaching a goal as quickly as possible. While our Mountain Car experiments show that FQI-log is able to achieve the goal of reaching the top of the hill before the episode terminates, FQI-log does not attempt to minimize the steps needed to reach the goal. We leave the issue of reaching goals quickly as a promising direction for future research.
Acknowledgement
We would like to thank Roshan Shariff for pointing out a bug in an earlier version of our proof.
Appendix
@@ Appendix A Preliminary results
In this section we introduce and prove some elementary inequalities connecting various metrics defined for nonnegative valued functions and a concentration result for the log-loss estimator taken from Foster and Krishnamurthy [2021], The concentration result gives a high probability upper bound on the error of the log-loss estimator, where the error is measured with the integrated binary Hellinger loss (to be defined below). We will base our analysis on this result. The elementary inequalities connect the integrated binary Hellinger loss to both the Hellinger distance and the so-called triangular discrimination metric, and they will allow us to reduce the analysis of our algorithm to studying the approximation error of the value functions constructed by our algorithm.
The concentration The analysis of the algorithm revolves around using the Hellinger distance, which is a distance between nonnegative valued functions. In particular, for $f,g\geq 0$ with measurable domain $\mathcal{X}$ and a measure $\lambda$ over $\mathcal{X}$, the Hellinger distance between $f$ and $g$ is defined as
$\operatorname{H}(f,g)=\frac{1}{2}\|\sqrt{f}-\sqrt{g}\|_{2,\lambda}\,.$
As it turns out, a version of the Hellinger distance the log-loss estimator is best
@@ A.1 Some basic inequalities
We start by defining what we call the binary Hellinger loss between two real numbers in the $[0,1]$ interval. size=,color=red!20!white,]Cs: so maybe this should have been $\ell_{\operatorname{H}}(p,q)$. In particular, we let $p,q\in[0,1]$ reals, let
$\operatorname{h}^{2}(p\,\|\,q):=\frac{1}{2}\left(\sqrt{p}-\sqrt{q}\right)^{2}+\frac{1}{2}\left(\sqrt{1-p}-\sqrt{1-q}\right)^{2}\,.$ (6)
Note that the Hellinger distance between two distributions $P$ and $Q$ over a common domain is defined as $\frac{1}{\sqrt{2}}\left\|\sqrt{p}-\sqrt{q}\right\|_{2,\lambda}$, where $p=dP/d\lambda$ and $q=dQ/d\lambda$ with $\lambda$ a distribution that is dominating both $P$ and $Q$ (thus, above, $p$ ($q)$ is the density of $P$ ($Q$) relative to $\lambda$, respectively). size=,color=red!20!white,]Cs: citation With this we can note that the binary Hellinger loss between $p$ and $q$ is the Hellinger distance between the Bernoulli distributions with means $p$ and $q$. Note that $0\leq\operatorname{h}^{2}(p\mathrel{\|}q)\leq 1$.
@@
For the first inequality in the statement, notice that
$\frac{(p-q)^{2}}{4(p+q)}\leq\frac{(p-q)^{2}}{2(\sqrt{p}+\sqrt{q})^{2}}=\frac{1}{2}\left(\sqrt{p}-\sqrt{q}\right)^{2}\,,$ (8)
where the inequality uses $(\sqrt{p}+\sqrt{q})^{2}\leq 2p+2q$. For the second inequality in the statement, we have
$\frac{1}{2}\left(\sqrt{p}-\sqrt{q}\right)^{2}\leq\frac{1}{2}\left(\sqrt{p}-\sqrt{q}\right)^{2}+\frac{1}{2}\left(\sqrt{1-p}-\sqrt{1-q}\right)^{2}=\operatorname{h}^{2}(p\,\|\,q)\,.\qed$
To state the next result, which is a simple corollary of the above inequality, we extend the definition of $\operatorname{h}^{2}$ so that it can be applied to function $f,g:\mathcal{X}\to[0,1]$. In particular, we let $\operatorname{h}^{2}(f,g):\mathcal{X}\to[0,1]$ defined via size=,color=red!20!white,]Cs: so maybe this should have been $\ell_{\operatorname{H}}(f,g,\cdot)$. Oh well..
$(\operatorname{h}^{2}(f,g))(x)=\operatorname{h}^{2}(f(x)\mathrel{\|}g(x))\,\qquad x\in\mathcal{X}\,.$
With this, a simple corollary of the first statement of this section is as follows:
@@
$\left\|\frac{f-g}{\sqrt{f+g}}\right\|_{2,\nu}\leq\sqrt{2}\left\|\sqrt{f}-\sqrt{g}\right\|_{2,\nu}\leq 2\left\|\operatorname{h}^{2}(f,g)\right\|_{1,\nu}^{1/2}\,.$
where for $f(x)=g(x)=0$, we define $\frac{f(x)-g(x)}{\sqrt{f(x)+g(x)}}=0$.
We call the quantity on the right-hand side as the square root of the integrated binary Hellinger loss between $f$ and $g$. Squaring all quantities above, we get the equivalent inequalities
@@
With words, this states, that the integrated binary Hellinger loss between $f$ and $g$ is lower bounded by their squared Hellinger distance, which is lower bounded, up to a constant factor, by the so-called triangular discrimination distance them. The latter is in essence the squared distance between $f$ and $g$, where their pointwise differences is rescaled with the square root of the sum between the functions. Because $|a-b|/\sqrt{a+b}\leq\varepsilon$ implies $|a-b|\leq\varepsilon\sqrt{a+b}\leq\varepsilon\max(\sqrt{a},\sqrt{b})$, we see that a bound on the rescaled distance between two values tightens the bound between them when both values are close to zero. In essence, this is what will take advantage of later in our analysis.
Proof.
Apply Lemma A.1 pointwise and then integrate both sides of the inequality over $\nu$, and then take the square root and simplify the constants. ∎

### v5

@@ 6.3 Asterix and Seaquest
As described earlier, we evaluate the deep RL variants of FQI-log and FQI-sq on the Atari 2600 games Asterix and Seaquest (Bellemare et al., 2013), and use the distributional RL algorithm C51 as an additional baseline. We adopt the data and experimental setup of Agarwal et al. (2020). The data consist of five batch datasets for each game, which were collected from independent training runs of a DQN learner (Mnih et al., 2015). Specifically, each batch dataset contains every fourth frame from 200 million frames of training; a frame skip of four and sticky actions (Machado et al., 2018) were used, whereby all actions were repeated four times consecutively and a learner randomly repeated its previous action with probability 0.25.
When the function class $\mathcal{F}$ of FQI-sq is given by a deep neural network, the algorithm is called DQN. To adapt FQI-log to the deep RL setting, we must switch the training loss from $\ell_{\text{sq}}$ to $\ell_{\log}$, and add a sigmoid activation layer to squash the output range to $[0,1]$. We henceforth refer to these algorithms as DQN-sq and DQN-log, respectively.
@@
    C51 replaces the Bellman optimality operator $\mathcal{T}$ with a modified “distributional Bellman operator”.
For our experiments we clip the targets of DQN-log by setting $v_{\min}=0$ and $v_{\max}=10$. Since the sigmoid activation of DQN-log is a specialization of the softmax activation to the binary case, DQN-log implements the changes S.3, S.2, and S.4 to the standard form of DQN-sq. Clipping the targets introduces a bias which we correct for by similarly clipping the targets of DQN-sq. Thus our benchmark results for DQN-sq include the change S.4. Clipping the targets of DQN-sq is novel to this work and improves performance, yielding a stronger baseline. We include a comparison with the traditional unclipped version of DQN-sq in Appendix C.
In our implementations of DQN-log and DQN-sq, we use the same hyperparameters reported by Agarwal et al. (2020). Figure 3 shows the undiscounted return as a function of the number of training epochs. On Seaquest, DQN-log outperforms DQN-sq and matches the performance of C51. In Asterix, DQN-log performs similarly to DQN-sq and both get lower return than C51. Overall, our results are inconclusive in this setting in regards to whether switching to log-losses suffices to reproduce the success of C51. However, the experiments confirm that switching to log-loss can be beneficial, as compared to using the squared loss and sometimes this switch alone is sufficient to compete with the more complex C51 algorithm.
Figure 3: Learning curves on Asterix and Seaquest. The results are averaged over $5$ datasets. The shaded regions represent one standard error of the mean. One epoch contains 100k updates.
@@ 7 Related Works
Wang et al. (2023) obtain small-cost bounds for finite-horizon batch RL problems under the distributional Bellman completeness assumption, which is more restrictive than our analogue, Assumption 3.4. Wang et al. (2024) refines the bounds of Wang et al. (2023), showing second-order bounds (which depends on the variance) for the same algorithm. They attribute their small-cost bound to the use of the distributional Bellman operator (Bellemare et al., 2023). However, their proof techniques only make use of pessimism (Buckman et al., 2021; Jin et al., 2021) and log-loss in achieving their small-cost bound. The use of pessimism is necessary for their proof in order to control the errors accumulated by use of the distribution Bellman operator during value iteration. We improve upon their work by proposing an efficient algorithm for batch RL that enjoys a similar small-cost bound without using the distributional Bellman operator or pessimism under a weaker completeness assumption.
Jin et al. (2020a) and Wagenmaker et al. (2022) obtain regret bounds that scale with the value of the optimal policy. However in their setting, the goal is to maximize reward. Therefore, their bounds only improve upon previous bounds (e.g., Azar et al. 2017; Yang & Wang 2019; Jin et al. 2020b) when the optimal policy accumulates very little reward. These bounds are somewhat vacuous as they only imply that regret is low when the value of the initial policy is already close to the value of the optimal policy, both of which are close to zero. Small-cost bounds give the same rates as small-return bounds, however, they are more attractive as the cost of the initial policy can be high while the cost of the optimal policy can be low. Small-cost bounds for online learning were given by Lee et al. (2020) for learning in tabular MDPs, while Kakade et al. (2020) derived such results for linear quadratic regulators (LQR). To our knowledge we give the first small-cost bound for batch RL under completeness (Antos et al., 2007; Munos & Szepesvári, 2008).
@@
Several works get small-cost bounds in contextual bandits (Allen-Zhu et al., 2018; Foster & Krishnamurthy, 2021; Olkhovskaya et al., 2023). In their Theorem 3, Foster & Krishnamurthy (2021) show that for batch contextual bandits, playing the greedy policy with respect to a reward function estimated via log-loss enjoys a small-cost bound, whereas in their Theorem 2 they show that if a reward function is estimated via squared loss, the greedy policy fails to achieve a small-cost guarantee. Our main result, Theorem 5.1 can be viewed as the sequel to their Theorem 3. Abeille et al. (2021) show that for stochastic logistic bandits, where the costs/rewards are Bernoulli, the regret can be made to scale with the variance of the optimal arm. This is simultaneously a small-cost and small-return bound. They achieve this bound by employing optimism together with maximum likelihood estimation (MLE). Janz et al. (2024) extend this result and show that this result continues to hold if the reward distribution comes from a “self-concordant”, single-parameter family.
@@
The theory literature on batch RL has largely focused on proving sample efficiency rates. Chen & Jiang (2019) proved that FQI-sq gets a rate optimal bound of $1/\sqrt{n}$ when realizability, concentrability and completeness hold. Foster et al. (2021) then show that if one assumes concentrability then completeness is a necessary assumption for sample efficient batch RL. Xie & Jiang (2021) prove that if one uses the stronger notation of concentrability from Munos (2003) then sample efficient batch RL is possible even with only realizability. To our best knowledge, the previous theoretical works on batch RL have only considered algorithms where value estimation used squared loss (e.g., Antos et al., 2007; Farahmand, 2011; Pires & Szepesvári, 2012; Chen & Jiang, 2019; Xie & Jiang, 2021).
Concurrent empirical work
The concurrent and independent empirical work of Farebrother et al. (2024) also advocates for log-loss, but they end up with the approach that is used in distributional RL, which reduces regression to multiclass classification. This is not only more complicated than using log-loss (more parameters), but also introduces irreducible bias, whereas our approach avoids this.
@@ 8 Conclusions
By proving that FQI-log is more sample efficient in MDPs with a small optimal cost $\bar{v}^{\star}$ than FQI-sq, we showed that in batch RL the loss function genuinely matters. We believe our result holds generally and can be extended to any batch RL setting where the squared loss has bounded error, such as when pessimistic methods are used. Another intriguing extension would be deriving small-cost bounds in batch RL with only realizability (and a stronger concentrability assumption), perhaps following the analysis of Xie & Jiang (2021). Wagenmaker et al. (2022) get small-return bounds for online RL in linear MDPs. Can we get small-cost bounds in online RL with linear function approximation? When and how? Finally, our mountain car experiments indicate that log-loss might perform well in goal-oriented MDPs (Bertsekas, 1995), where the learner is tasked with reaching some goal and this is possible. However, in this formulation there is no “pressure” for reaching the goal as quickly as possible. It remains to be seen how this could be incorporated in algorithms like ours. (Staying away from failure zones for as long as possible can be easily formulated with the help of discounting.) Our experiments concerning whether switching to the log-loss is sufficient to replicate successes of the more complex C51 distributional RL algorithm were inconclusive. Here, it will be interesting to investigate whether switching to log-loss, but also keeping the losses used by C51 as auxiliary loss can lead to a method that outperforms both C51 and our method.
Acknowledgments
The authors would like to thank Roshan Shariff for pointing out a bug in an earlier version of our proof. Csaba Szepesvári also gratefully acknowledges funding from the Canada CIFAR AI Chairs Program, Amii and NSERC.
Impact Statement
This paper presents work whose goal is to advance the field of reinforcement learning theory. There are many potential societal consequences of our work.
@@ Appendix A Preliminary results
In this section we introduce and prove some elementary inequalities that connect useful metrics on function spaces, and then state a concentration result of Foster & Krishnamurthy (2021) for the log-loss estimator. The concentration result gives a high probability upper bound on the error of the log-loss estimator, as measured by the integrated binary Hellinger loss (defined below). This result is central to our analysis. The elementary inequalities connect the integrated binary Hellinger loss to both the Hellinger distance and the triangular discrimination, and will reduce the analysis of our algorithm to studying the approximation error of its value function estimates $\{f_{j}\}_{j=1}^{k}$.
The analysis of FQI-log revolves around controlling the Hellinger distance, which is a distance between nonnegative integrable functions. In particular, for $\lambda$-integrable functions $f,g\geq 0$, the Hellinger distance between $f$ and $g$ is defined as
$\operatorname{H}(f,g)=\frac{1}{\sqrt{2}}\left\|\sqrt{f}-\sqrt{g}\right\|_{2,\lambda}\,.$
@@ A.1 Some basic inequalities
Given real numbers $p,q\in[0,1]$, we define the binary Hellinger loss of $p$ and $q$ as
$\operatorname{h}^{2}(p,q)=\frac{1}{2}\left(\sqrt{p}-\sqrt{q}\right)^{2}+\frac{1}{2}\left(\sqrt{1-p}-\sqrt{1-q}\right)^{2}\,,$ (7)
and immediately observe that $0\leq\operatorname{h}^{2}(p,q)\leq 1$. Note that the Hellinger distance between two distributions $P$ and $Q$ over a common domain is defined as $\frac{1}{\sqrt{2}}\left\|\sqrt{p}-\sqrt{q}\right\|_{2,\lambda}$, where $p=dP/d\lambda$ and $q=dQ/d\lambda$ are the densities of $P$ and $Q$ with respect to a dominating distribution $\lambda$. size=,color=red!20!white,]Cs: citation Thus the binary Hellinger loss between $p$ and $q$, $\operatorname{h}^{2}(p,q)$, is the squared Hellinger distance between Bernoulli distributions with means $p$ and $q$.
@@
$\frac{1}{4}\frac{(p-q)^{2}}{p+q}\leq\frac{1}{2}\left(\sqrt{p}-\sqrt{q}\right)^{2}\leq\operatorname{h}^{2}(p,q)\,,$ (8)
@@
If $p=q=0$ then equality holds trivially, and otherwise $(\sqrt{p}+\sqrt{q})^{2}\leq 2(p+q)$ implies
$\frac{(p-q)^{2}}{4(p+q)}\leq\frac{(p-q)^{2}}{2(\sqrt{p}+\sqrt{q})^{2}}=\frac{1}{2}\left(\sqrt{p}-\sqrt{q}\right)^{2}\leq\frac{1}{2}\left(\sqrt{p}-\sqrt{q}\right)^{2}+\frac{1}{2}\left(\sqrt{1-p}-\sqrt{1-q}\right)^{2}=\operatorname{h}^{2}(p,q)\,.\qed$
The next result holds for an extended definition of $\operatorname{h}^{2}$ that replaces the inputs $p,q\in[0,1]$ with functions $f,g:\mathcal{X}\to[0,1]$. Given such functions, we define $\operatorname{h}^{2}(f,g):\mathcal{X}\to[0,1]$ by
$(\operatorname{h}^{2}(f,g))(x)=\operatorname{h}^{2}(f(x),g(x))\,,\qquad x\in\mathcal{X}\,.$
With this definition in hand, the following is a straightforward corollary of Lemma A.1.
@@
$\left\|\frac{f-g}{\sqrt{f+g}}\right\|_{2,\nu}\leq\sqrt{2}\left\|\sqrt{f}-\sqrt{g}\right\|_{2,\nu}\leq 2\left\|\operatorname{h}^{2}(f,g)\right\|_{1,\nu}^{1/2}\,,$
where for $f(x)=g(x)=0$ we define $\frac{f(x)-g(x)}{\sqrt{f(x)+g(x)}}=0$.
We call the quantity $\left\|\operatorname{h}^{2}(f,g)\right\|_{1,\nu}$, which appears on the right hand side above, the integrated binary Hellinger loss between $f$ and $g$. size=,color=orange!20!white,]SR: this name will never be mentioned again Squaring all quantities and dividing through by 4 yields the equivalent inequalities
@@
In words, the integrated binary Hellinger loss between $f$ and $g$ is lower bounded by their squared Hellinger distance, which itself is lower bounded by one quarter of the triangular discrimination between them. The latter is essentially the squared distance between $f$ and $g$, but rescaled pointwise by $1/\sqrt{f+g}$. Because $|a-b|/\sqrt{a+b}\leq\varepsilon$ implies $|a-b|\leq\varepsilon\sqrt{a+b}$, we see that a bound on the rescaled distance between two values tightens the bound between them whenever $\sqrt{a+b}<1$. Exploiting this property is key to our later analysis.
Proof of Corollary A.2.
Apply Lemma A.1 pointwise and then integrate both sides of the inequalities over $\nu$, before taking square roots and multiplying by 2 to simplify the constants. ∎